OpenStudy (anonymous):

Find the number of years it would take for $1200 to earn simple interest of $324 at an annual interest rate of 6% per year.

OpenStudy (mathstudent55):

I = prt I = interest p = principal amount r = annual interest rate t = time in years

OpenStudy (anonymous):

give me the awnser please

OpenStudy (mathstudent55):

$324 = $1200 * 0.06 * t $324 = $72t t = 324/72 t = 4.5